OpenStudy (anonymous):

(x^2+4x)-2x=x^2+?x WHAT IS THE MISSING COEFFICIENT

OpenStudy (solomonzelman):

(x^2+4x)-2x ---> x^2+2x x^2+2x = x^2+?x the ?=...

OpenStudy (anonymous):

I honestly have no clue how too do any of this. I've missed school the past week.

OpenStudy (solomonzelman):

you can subtract x^2 from both sides.

OpenStudy (anonymous):

Both sides of what

OpenStudy (anonymous):

Thanks I get it now

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

Your not supposed too give the answer where mods can see it. lol . ^^^

OpenStudy (anonymous):

Oh.. sorry